### v1

# Accelerating Tactile Internet with QUIC: An Exploration of its Security and Privacy Attacks

Abstract
The Tactile Internet paradigm is set to revolutionize human society by enabling skill-set delivery and haptic communication over ultra-reliable, low-latency networks. The emerging sixth-generation (6G) mobile communication systems are envisioned to underpin this Tactile Internet ecosystem at the network edge by providing ubiquitous global connectivity. However, apart from a multitude of opportunities of the Tactile Internet, security and privacy challenges emerge at the forefront. We believe that the recently standardized QUIC protocol, characterized by end-to-end encryption and reduced round-trip delay would serve as the backbone of Tactile Internet. In this article, we envision a futuristic scenario where a QUIC-enabled network uses the underlying 6G communication infrastructure to achieve the requirements for Tactile Internet. Interestingly this requires a deeper investigation of a wide range of security and privacy challenges in QUIC, that need to be mitigated for its adoption in Tactile Internet. Henceforth, this article reviews the existing security and privacy attacks in QUIC and their implication on users. Followed by that, we discuss state-of-the-art attack mitigation strategies and investigate some of their drawbacks with possible directions for future work.
I Introduction
The Tactile Internet [1], driven by recent technological advancements, focuses on transmitting touch in real-time via suitable haptic equipment and robotics over a robust communication network. Tactile Internet promises to deliver real-time control through ultra-responsive and ultra-reliable network connectivity over a distance to create a more sustainable ecosystem. This is expected to revolutionize the entire human society by shifting the focus from mere content delivery to a new system of skill-set delivery over the Internet. However, its implementation at the infrastructure level demands ensuring low latency with highly reliable communication and privacy coupled with a very high level of security. This is because Tactile Internet is designed to serve highly critical applications where human senses (touch, video, and audio) and muscular movements directly interact with machines, i.e., robots. Hence, cybersecurity and quick reaction time ($<1$ ms) are two primary requirements of the Tactile Internet [1].
Any significant delay can induce cyber-sickness in users due to unsynchronized senses, occurring as a result of a time lag ($>1$ ms) between visual and tactile movements. Achieving global connectivity with a high data rate and minimizing jitter are additional challenges. Given such non-elastic requirements, the emerging sixth-generation (6G) communication system, offering ubiquitous connectivity, is poised to serve as the Tactile Internet’s network backbone. While existing 5G systems meet the throughput requirements, the integration of terahertz (THz) communication and millimeter-wave (mmWave) in 6G ensures higher data rates and ultra-low-latency (0.1 ms delay), making it ideal for the Tactile Internet [2].
Figure 1: A high-level architectural view of a QUIC-enabled futuristic Tactile Internet infrastructure using 6G as the underlying mobile communication network.
It is evident from the above discussion that security/privacy breaches and packet losses in the Tactile Internet environment can lead to detrimental effects, particularly if it is used in mission-critical applications. In this regard, the traditional transport layer protocols such as TCP and UDP are inadequate as they are not designed to ensure security. Moreover, the “TCP head-of-line blocking” problem  [3] poses a significant bottleneck for Tactile Internet applications, as packet loss from one source leads to buffering of successive packets from all sources (e.g., different sensors). Given the high redundancy in sensory data, receiving every generated packet is not mandatory. Hence, the Internet Engineering Task Force’s (IETF) recently standardized transport protocol, QUIC (see: §II), is a suitable alternative.
QUIC is a connection-oriented, end-to-end encrypted protocol based on UDP and built on top of TLS 1.3, thus supporting $0-RTT$. QUIC allows sending parallel and independent data streams which are logically separate from one another. This ensures fast and reliable in-order delivery of independent data streams, thereby avoiding the head-of-line blocking problem faced by TCP [3]. Though QUIC was primarily designed to handle Web traffic, its unique features (low-latency delivery, security, $0-RTT$ performance benefits) are suitable for interactive Tactile Internet applications that need to transmit kinesthetic and tactile data collected from various sensors with potential low delay and high security. Moreover, QUIC is already under the deployment stage across multiple application scenarios such as DNS, VPNs and beyond other than just the Web [3]. Inspired by these facts, we envision QUIC as a potential candidate for 6G-based Tactile Internet applications as depicted in Fig. 1. Consequently, a deeper investigation of the various security (see: §II) and privacy (see: §III) challenges in QUIC is needed.
This article provides a comprehensive overview of the security and privacy issues in QUIC. We believe this will help practitioners and system designers take appropriate measures before deploying QUIC in a Tactile Internet environment. We also discuss a few mitigation strategies and highlight some of their drawbacks (see: §IV), providing guidance for the research community to delve deeper and propose alternative or improved solutions to the challenges outlined in this article.
Figure 2: QUIC connection establishment with attack vectors. The packets or tuples under attack are represented in blue, whereas the attacks are highlighted in red.
II Security Attacks on QUIC
We begin this section by providing a deep dive into QUIC handshakes, as several security attacks exploit the packets exchanged during QUIC connection establishment. We then discuss the relevant security attacks on QUIC (see: Table I) applicable to a Tactile Internet environment.
QUIC employs end-to-end encryption, necessitating agreement on multiple parameters between endpoints (client and server). This agreement occurs through a TLS 1.3 handshake, as depicted in Fig. 2. A typical QUIC segment’s header fields are also illustrated. When a client lacks prior knowledge of a server, it initiates the process by sending an Initial Client Hello (CHLO) message to the server. Upon receiving CHLO, the server responds with a reject reject (REJ) packet to the client, containing tuples: (i) server config (scfg) with Diffie-Hellman public value, (ii) source-address token (stk) comprising timestamp from the server and client’s IP address, (iii) server authentication certificate chain, and (iv) server config signature. After receiving the server config, the client authenticates it using the certificate chain and the server signature. The client then creates a Complete CHLO, including its ephemeral Diffie-Hellman key, and sends it back to the server. With the completion of the handshake, the client obtains the initial keys for communication, enabling it to send an encrypted connection request to the server. The server responds with a server hello (SHLO) packet, encrypted with initial keys and the server’s ephemeral keys. Both client and server then calculate their forward secure keys to be used for subsequent communication rounds. The client can now engage in a 00-RTT handshake, bypassing the initial handshake mechanism, by sending a Complete CHLO.
II-A 0-RTT Attack
During the initiation of a connection setup between the client and server, a 00-RTT attack can occur when the attacker intercepts the unencrypted initial CHLO packet and spoofs it (Fig. 2). This opens the door to two potential attacks [4]:
QUIC RST Attack: Attacker sends a public reset packet to the client, deceiving it into thinking that the server has rejected the connection. This leads the client to proactively abandon the connection.
Version Forgery Attack: The attacker poses as a server and transmits a version negotiation packet to the client, containing a version unsupported by the client. This prompts the client to either downgrade or abandon the connection.
In this regard, Cao et al. [4] have detailed the core concepts and implementation of the 00-RTT attack, emphasizing its potential to result in denial-of-service. They have introduced a mathematical attack description model based on finite state machines to demonstrate the QUIC protocol’s vulnerability, showcasing the attack process and confirming QUIC’s susceptibility to the described attacks.
II-B Replay Attack
A replay attack occurs when an intruder intercepts a legitimate network transmission and later resends it to deceive the system into treating the re-transmitted data as authentic. Lychev et al. [5] conducted two types of replay attacks (Fig. 2) on the QUIC implementation in Chromium:
Server Config Replay Attack: This occurs when an adversary replays the server’s public value (scfg) to clients who have sent an initial connection request to the server, without being detected by the server. Consequently, these clients generate the initial key and send it to the server, which on being unable to verify their identity, rejects the packets. While the secrecy of the entities involved remains largely unaffected, this attack results in unnecessary consumption of computational resources.
Source Address Token Replay Attack: In this attack, the intruder replays the source-address token (stk) of a client to the server that issued it, allowing the creation of multiple additional connections. This will make the server establish both initial keys and final forward-secure keys. Although subsequent handshake steps would fail, the attacker could potentially launch a denial-of-service (DoS) attack on the server by generating numerous connections on behalf of various clients, leading to a depletion of computational and memory resources.
Fischlin et al. [6] addressed replay attacks in the 00-RTT QUIC handshake. While 00-RTT uses registers for storing nonces to prevent repetition, in QUIC, once the server rejects the 00-RTT message, the attacker resends the rejected message using a second key to ensure proper delivery. This results in duplicate processing of the same data by the server.
TABLE I: Overview of various Security and Privacy Attacks and their Effects on QUIC
Category Attack Surface Attack Name Effects Reference
Security Attack Initial CHLO message 0-RTT Attack
Downgrade; Client-side
Connection Rejection
[4]
REJ message Replay Attack
Connection Failure;
Server DoS
[5, 6]
Source-Address Token (stk)
Manipulation Attack
Connection Failure; Server Load [5]
Packet Header
Connection ID (cid)
Manipulation Attack
[5]
Request Forgery Attack Traffic Amplification [7]
Handshake and
Packet Header
State-Overflow and
Resource Exhaustion Attacks
Fake connections;
Resource Wastage
[8]
QUIC Server QUIC Flood DDoS Attack
Resource Exhaustion;
System breakdown
[8, 9]
Privacy Attack Traffic traces and Communication Channel User Tracking via QUIC
User Profiling;
Targetted Advertising
[10]
Website Fingerprinting User Privacy breach [11, 12]
II-C Manipulation Attack
A manipulation attack aims to compromise the key agreement process, leading the client and server to agree on distinct keys. It is achieved by altering unprotected packet fields, such as connection id (cid) or the source-address token (stk) (Fig. 2) that are utilized as input to the key derivation process. Lychev et al. [5] describe the following two types of manipulation attacks in QUIC:
Connection ID Manipulation Attack: Occurs when an attacker generates a fresh cid, causing the server and client to perceive different cid values. Though the handshake begins normally, the server and the client have distinct encryption keys due to different cid values. Consequently, decryption fails, and failed packets are buffered until the handshake is completed. In due time, this situation leads to the disconnection of the link. However, the error message gets encoded using the original encryption key, thereby preventing its decryption and preserving the state until it times out.
Source-Address Token Manipulation Attack: In this attack, the attacker generates separate stk values for both the client and server, resembling a CID manipulation attack. The handshake typically begins with the transmission of CHLO messages. However, since encryption keys rely on stk as an input, the client and server generate distinct encryption keys, leading to decryption failure. As a result, the client buffers failed CHLO messages and retransmits them for 10 seconds. After a timeout, the client sends an encrypted error message to the server. Yet, as the server possesses initial encryption keys, the connection remains intact, and this state may persist for 10 minutes.
II-D Request Forgery Attack
This attack involves instigating a client to send deceptive requests to another client. In [7], Gbur et al. assume a scenario where the attacker has complete control over packets delivered to the victim, enabling three types of forgery attacks:
Server Initial Request Forgery (SIRF): The attacker initiates a QUIC handshake with the server, where the packet’s source IP address and port fields are spoofed. This tricks the victim to assume that the connection is from a genuine server.
Version Negotiation Request Forgery (VNRF): In this attack, the attacker sends an unknown version in the client’s packet, prompting the QUIC server to respond with a version negotiation packet. This action triggers the version negotiation functionality.
Connection Migration Request Forgery (CMRF): While QUIC’s connection migration feature offers advantages, a drawback exists: the server is unable to differentiate between a genuinely migrated client address and a spoofed one. In this scenario, an attacker conducts a legitimate handshake with the server to establish a new connection, then spoofs an arbitrary packet to obtain the source address and sends it to the server. Consequently, the server, upon detecting a new source address, unwittingly establishes the connection and transmits UDP packets to the manipulated address.
II-E State-Overflow and Resource Exhaustion Attacks
In these attacks, adversaries masquerade as legitimate clients and initiate full handshakes with QUIC servers. The server responds with a unique Source Connection ID (SCID) and the corresponding TLS certificate, thus allocating resources to maintain connection states. To overwhelm the server, the adversarial client randomly spoofs source ports and IP addresses, flooding the server with multiple handshake requests. This results in the creation of numerous unique SCIDs and concurrent connection states, leading to server overload, state overflow, and resource exhaustion — similar to TCP SYN floods, where legitimate client requests may also be rejected. The QUIC design’s primary vulnerability lies in the lack of client verification during the initial round-trip of the QUIC full handshake.
A study [8] evaluates these attacks on real-world QUIC traffic based on active measurements using the UCSD network telescope. Results show that, similar to TCP SYN floods, the QUIC handshake is susceptible to resource exhaustion attacks. Notably, 98% of these attacks target QUIC servers of well-known companies, with Google being the primary target (58%) and Facebook accounting for 25% of the attacks. While QUIC originally supports RETRY messages to mitigate resource exhaustion attacks, their implementation adds an extra round-trip time (RTT), conflicting with QUIC’s claimed performance gains. This happens as a RETRY message typically precedes a QUIC handshake and forces the client to prove its authenticity by responding with a unique token. The study [8] emphasizes that during experimentation, no RETRY messages were observed, indicating a lack of practical defense mechanisms in QUIC deployments.
II-F QUIC Flood DDoS Attack
As previously explained, a QUIC flood DDoS attack occurs when an attacker overwhelms a QUIC server with a large volume of data, causing the victim server to slow down significantly or crash. Defending against DDoS attacks in QUIC is challenging because QUIC is based on UDP, which provides minimal or no information for blocking illegitimate traffic. Additionally, as QUIC packets are encrypted, the victim server cannot easily verify the legitimacy of the data source. While some Content Distribution Networks (CDNs) like Cloudflare have been successful in mitigating QUIC floods [9], a study in [8] indicates that these attacks persist in the real world. In [8], the authors reported that the Internet is vulnerable to four QUIC flood attacks every hour. Among these, 51% of the attacks occur concurrently with TCP/ICMP floods, while another 40% target the victim sequentially. However, QUIC floods have a shorter duration, lasting only $255\ seconds$, compared to their TCP/ICMP counterparts. However, such shorter duration may lead to faster resource exhaustion or system breakdown, presenting an aspect that is yet to be explored.
III Privacy Attacks on QUIC
This section reviews two specific privacy threats that are posed to QUIC-based user applications.
III-A User Tracking via QUIC
Online tracking threatens user privacy by exploiting users’ browsing habits, which may reveal sensitive information. It can then be used for profiling, web analytics, and targeted ads. To counter this, browsers need robust privacy protection. While QUIC, being encrypted is widely adopted across Web browsers as it aims to ensure user privacy, it remains susceptible to tracking through the following two mechanisms:
Linking several website visits by the same user: The source-address-token is a unique data block included in the reject (REJ) message sent by the server to a client during initial connection setup. The client caches it and presents to the server for every new 00-RTT connection setup with the same server. This enables the server to identify subsequent connection requests from the same user which enables linking different website visits to the same usernames.
User tracking across multiple sessions: QUIC clients cache server-config (part of REJ message), containing a unique 16-byte server config identifier (SCID) assigned by the server to each user. This SCID allows Web servers and potential attackers to link the initial connection request to subsequent requests using the same SCID within CHLO messages.
The study [10] reveals QUIC’s privacy violation, enabling trackers to map users based on tokens provided during a 00-RTT connection setup attempt. Analysis of popular browsers (Chrome) revealed insufficient protective measures against Web user tracking. Such monitoring was particularly effective in resource-constrained scenarios due to lower bandwidth requirements and delays compared to HTTP cookies or traditional browser fingerprinting.
In a use-case scenario (Fig. 3), despite using encrypted QUIC connection, an attacker observing traffic traces during a real-time haptic signal transfer session (such as remote music training), can still gather information about students and the teacher. This information could be strategically exploited for user profiling and collecting sensitive details like gender and location.
Figure 3: A use-case scenario of Online Education where a network attacker can eavesdrop on the QUIC connection to launch privacy attacks.
III-B Website Fingerprinting
An encrypted transmission protocol is susceptible to website fingerprinting (WFP) attacks if adversaries can deduce a user’s visited websites by monitoring the transmission channel. As QUIC encrypts data, adversaries target unencrypted handshake packets. Further, as QUIC is based on a request-response model, it allows attackers to distinguish traffic from different Web resources due to the relatively fixed browser rendering sequence [11]. This weakens privacy benefits achieved through encryption under WFP attacks.
Zhan et al. [11] explored WFP attack vulnerabilities in both QUIC variants: (i) GQUIC (Google’s QUIC) and (ii) IQUIC (IETF’s QUIC). GQUIC is more vulnerable in restricted traffic scenarios, while both protocols exhibit similar vulnerabilities under typical traffic. Multiple WFP attacks with only 40 packets achieved 95.4% and 95.5% accuracy for GQUIC and IQUIC, respectively.
To counter WFP attacks, IQUIC introduces a PADDING frame, but studies suggest network-level padding inefficiency against adversaries capable of observing traffic traces. Current research [12] focuses on integrating defenses directly into client applications (e.g., browsers) by effectively using the PADDING frame to thwart adversary’s attempt of traffic analysis in Virtual Private Networks (VPNs).
A summary of various security and privacy attacks on QUIC is presented in Table I.
IV Potential Mitigation Strategies
In QUIC, connection establishment between the endpoints is achieved through a TLS 1.3 handshake where the client sends an initial ClientHello to the server over a UDP socket. However, this introduces several security and privacy challenges as sensitive parameters are negotiated between the endpoints during such TLS connections. Though TLS 1.3 encrypts most of the information exchanged, an on-path attacker can still learn certain sensitive private information such as Server Name Identification (SNI) extension (see: § II) available in the ClientHello message. It is also evident from § II that most of the attacks occur due to the unencrypted packet exchange during connection establishment as eavesdroppers get access to a plethora of metadata during this phase.
This motivated the IETF to explore and encrypt the handshake messages apart from just the application data. In a work-in-progress document [^1] https://datatracker.ietf.org/doc/draft-ietf-tls-esni/16/, IETF proposes the inclusion of the Encrypted Client Hello (ECH) extension into the TLS to mitigate the security risks and improve the privacy of the Internet as a whole. The goal of ECH is to protect all sensitive handshake parameters such as preshared key, server name, etc. by encrypting the entire ClientHello, thereby closing the gap left by TLS 1.3 and Encrypted Server Name Identification (ESNI), the predecessor of ECH.
An extensive study was performed by Bhargavan et al. [13] to prove that TLS 1.3 with ECH indeed ensures basic security goals (like confidentiality, authenticity etc.), handshake privacy, and downgrade resistance. They provide a formal analysis of the privacy properties of TLS 1.3 in the presence of ECH using a symbolic protocol analyzer, named “ProVerif” to justify the security of ECH. However, the authors didn’t study privacy-preservation of ECH in the presence of traffic analysis. Therefore, any attack that relies on observing traffic patterns including the length of application messages and their timings would be something interesting to study in the future.
In another recent study [14], Trevisan et al. explore whether an attacker can recover the domain names of websites visited by a user in the presence of ECH by passively observing the traffic traces. The study reveals that simple machine learning models are capable of recovering a user’s visited websites with high accuracy. Their experiments reveal that ECH is only able to protect $\approx 7.5\%$ flows. In the end, they also evaluate the efficacy of standard padding-based techniques to thwart attacks on ECH. However, it is observed that the attackers are still able to gain a sufficiently large amount of information with just a minimal decrease in the effectiveness of the attacks launched. Hence, as a future work, it is important to improve ECH to increase its resistance to traffic analysis. Designing more robust techniques is necessary for offering better protection to end-users in terms of privacy.
While ECH can be a potential mitigation strategy against some of the security attacks (discussed in see: § II), Oblivious HTTP [^2] https://datatracker.ietf.org/doc/draft-ietf-ohai-ohttp/ and MASQUE (Multiplexed Application Substrate over QUIC Encryption) [^3] https://datatracker.ietf.org/doc/draft-ietf-masque-connect-ip/ are other potential work in progress technologies for the privacy attacks highlighted in § III. Private relay [15] offers privacy-preserving communication by establishing multiple proxy servers between the client and the server. This ensures that no attacker has visibility over both the user’s IP address and their access history. In this way, it separates the client’s IPs from servers and uses multi-hop MASQUE for proxying IP packets. On the contrary, Oblivious HTTP offers to send encrypted HTTP messages from a client to a gateway via a trusted relay service. This ensures that nobody sees the client’s IP address as well as the plaintext HTTP message remains hidden from the relay resources.
Overall, as a potential future work, researchers should consider designing privacy-preserving extensions or modifications to the QUIC protocol in order to mitigate potential privacy risks, such as by employing encryption for sensitive metadata and exploring differential privacy techniques. It is also crucial to integrate multi-factor authentication to fortify the trustworthiness of QUIC connections.
V Conclusion
The Tactile Internet is expected to bring a paradigm shift in all spheres of the connected society. However, this is only possible with tight integration of the network infrastructure, transport, and application (Tactile Internet) layers. Moreover, addressing the security and privacy challenges is crucial for the success of the Tactile Internet. In this article, we envisioned a QUIC-accelerated Tactile Internet that leverages 6G as the underlying network infrastructure for achieving low-latency, resilient global connectivity. We explored this ecosystem from a security and privacy perspective and presented several attacks that can make Tactile Internet vulnerable by targeting QUIC. To this end, we presented a few potential mitigation strategies that can be the future directions to plug the present security and privacy gap in this ecosystem, to help the Tactile Internet transition the connected society into a more sustainable future.